OpenStudy (anonymous):

Could somebody help me solve this please ? √13-N = √N-7

OpenStudy (anonymous):

@sourwing could you please help me ?

OpenStudy (solomonzelman):

what are the roots and what are not the roots?

OpenStudy (solomonzelman):

is 13-N and N-7 all included inside the the square roots ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \sf \sqrt{13-N}= \sqrt{N-7} }\) square both sides \(\Large\color{blue}{ \sf 13-N= N-7 }\) add N to both sides \(\Large\color{blue}{ \sf ~~~+N~~~~~+N }\) \(\Large\color{blue}{ \sf 13= 2N-7 }\) add 7 to both sides \(\Large\color{blue}{ \sf +7~~~+7 }\) \(\Large\color{blue}{ \sf 20= 2N }\) divide both sides by 2, ... \(\Large\color{blue}{ \sf 10= N }\)

OpenStudy (anonymous):

you're the bomb.com can i ask you another question? @SolomonZelman

OpenStudy (solomonzelman):

Thx for calling me that, although I am by far not the best mathematician in OS, closer to the worse one....

OpenStudy (solomonzelman):

Yeah, sure, go ahead and ask.

OpenStudy (anonymous):

√2x-18 = √x-5

OpenStudy (solomonzelman):

This time, I am going to tell you what to do, and you will do the steps I tell you for me. OK ?

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \sf \sqrt{2x-18} =\sqrt{x-5} }\) 1) Square both sides 2) subtract x from both sides 3) add 18 to both sides tell me the answer (I would prefer you to show your steps in here) .

OpenStudy (anonymous):

this might sound dumb but how do you square the sides ? is that when you multiply them by two ?

OpenStudy (solomonzelman):

I am going to show you .... once you know that \(\Large\color{blue}{ \bf \sqrt{a} =\sqrt{a} }\) you also know that \(\Large\color{blue}{ \bf (\sqrt{a})^{~2} =(\sqrt{a})^{~2} }\) and therefore you know that \(\Large\color{blue}{ \bf a =a }\) when you square both sides, you are simply raising each side to the second power, just like i did in the first problem.

OpenStudy (anonymous):

√2x-18^2=√x-5^2 like this??

OpenStudy (solomonzelman):

when you raise a square root of a number to a second power, the root cancels right?